OpenStudy (anonymous):

Please help Simplify: 9x+7/3 + 3x-8/3 I know the answer is 12x-1/3 would it be three or five. I'm confused.

OpenStudy (wwwhhhaaattt?):

\[12x-\frac{ 1 }{ 3 } OR 12x-\frac{ 1 }{ 5 }\]

OpenStudy (anonymous):

huh?

OpenStudy (wwwhhhaaattt?):

Is that what it looks like?

OpenStudy (wwwhhhaaattt?):

\[12x-\frac{ 1 }{ 3 }...OR...12x-\frac{ 1 }{ 5 }.....Is..this..what..\it..would..be?\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Thats what it looks like

OpenStudy (wwwhhhaaattt?):

Ok, Thats what it looks like?

OpenStudy (anonymous):

Yes

OpenStudy (wwwhhhaaattt?):

Ok, Hold on

OpenStudy (anonymous):

Okok

OpenStudy (anonymous):

6/7*4 IS THE ANSWER

OpenStudy (wwwhhhaaattt?):

\[\frac{ 4x+7 }{ 3 } + \frac{ 3x-8 }{ 3 }\]

OpenStudy (wwwhhhaaattt?):

@chordae I told you no more Direct Answers. So please Stop

OpenStudy (wwwhhhaaattt?):

Is that what it looks like?

OpenStudy (anonymous):

Nope.

OpenStudy (anonymous):

ill show you what I got.

OpenStudy (anonymous):

NO I LOVE GIVING THE ANSDWERS IT MAKES ME FEEL SMART

OpenStudy (wwwhhhaaattt?):

Brb